The revenue for a toaster company is directly proportional to the amount spent on advertising. When the owner spent $5,000 one month, the revenue was $600,000. If the revenue in the next month was $900,000, how much did the company spend on advertising?
I have come up with:
600000=5000k
K=120
900000=120x
X=7500
Is this correct????

Answer by checkley71(8403)

5,000/600,000=X/900,000
600,000X=5,000*900,000
600,000X=4,500,000,000
X=4,500,000,000/600,000
X=7,500
LOOKS GOOD TO ME.